Demand curve for an imperfect market is given as 100/(1+P)2 what price will the firm charge to maximize profit mc is equal to 2
1
Expert's answer
2020-04-08T09:38:26-0400

For imperfect competition MR=MC; MR=(P*Q)'.

From the demand equation P= (100/Q)1/2-1;

PQ'=(Q*((100/Q)1/2-1)'=5/Q1/2-1=2;

Q=(5/3)2; P=(100/(5/3)2)1/2=5

So, P=5